The first four cards could all be of different suits.
The next four cards might also all be of different suits.
This would mean that you would have two of each suit.
Then you might have to draw a ninth card to be sure of having three of the same suit.
What are the considerations for the choice of sample size?
Some considerations are:
What does it mean to say that the chances of getting a head when a coin is tossed are one-half?
It means that if the coin were tossed an infinite number of times,
half of the tosses would come up heads and half would be tails.
What is the probability of not rolling a sum of ten with two fair dice?
on a die of six 3\12 or if you want the fraction simplified 1/4
What are the odds if being born on a weekends?
The answer depends on a number of factors. The main one is whether, in the medical profession's judgement, there are risks associated with the pregnancy. If so, they might prefer a hospital birth. Since staffing levels are lower at weekends, many hospitals prefer to avoid weekend births and so the probability of being born on a weekday is increased.
What are the odds of getting a number graeater than 4 on a dice?
The chances of getting a number greater than 4 on a standard 6 face dice is a 2/6 chance which simplifies to a 1/3 chance. This is because their are only 2 numbers greater than 4.
If a random variable X has a Normal distribution with mean M and standard deviation S, then
Z = (X - M)/S
What is the probability of getting a sum of 7 and 11 when two number cubes are rolled?
Pr(7) = 1/36
Pr(11) = 2/36 = 1/18
What is the probabilty of a number lower than 25 being odd or a factor of 4?
The question is being answered assuming that by "number" you mean a positive integer. The probability that the number meets these requirements is 14/24 or 7/12.
If three number cubes are rolled all at once how many ways can you get 2 of the same number?
I'm going to assume you mean standard dice, all three distinguishable, and exactly two of the same number.
First break the problem into smaller parts. First, determine the odds of getting the same number on the first two dice, and a different number on the third. Then just multiply this result by 3 and 216 to account for the first and last dice being the same, or the last two dice, then compute the total number of possibilities.
P(two dice being the same and last being different) = P(second die being the same as the first) * P(third die being different than the other two) = (1/6)(5/6) = 5/36
3*216*5/36 = 90 possibilities out of 216
13
In a bag there 2 red 3yellow 4green and 6 blue marbles What is the probability of P Blue?
6 out of 15
If two fair dice were rolled what is the probability of them rolling a sum of 11?
x
1
2
3
4
5
6
1
2
3
4
5
6
7
2
3
4
5
6
7
8
3
4
5
6
7
8
9
4
5
6
7
8
9
10
5
6
7
8
9
10
11
6
7
8
9
10
11
12
Therfore, there is a 2 in 36 (or 1 in 18) chance of rolling a total of 11.
When tossing 2 fair 6 sided number cubes what is the probability of tossing a sum of ten?
The first cube can land in any one of 6 different ways.
The second cube can land in any one of 6 different ways.
Together, they can land in any one of (6 x 6) = 36 different ways.
Sum of ten can come up in 3 different ways:
4 - 6
6 - 4
5 - 5
The probability of tossing a ten is (3/36) = 1/12 = 8-1/3 %
Would it be possible for one to use a gigabyte in one months time with a mobile laptop?
It depends on what you are doing. If you download a lot of movies or music or take a lot of pictures, you can use a gigabyte in a few hours or a day. If you do nothing but type up word documents on your computer, you might not have over a gigabyte of typed documents in a year's time. Yes, it is possible to use a Gigabyte a month, it is also possible to use a gigabyte an hour or a gigabyte a year, it just depends on what you do on your computer and how you are using it.
Can parents be B pos and have a kid A B p?
No. If both parents are B+ then their children will be B+ (75%) or O+ (25%).
When determining the probability that two events happen at the same time, you convert the percents to decimals and then multiply the percents together. Therefore, 30 percent, or .3, times 50 percent, or .5
.3 x .5 = .15
Converting back into a percentage, the answer is 15% probability that you will get both. 10% is therefore incorrect.